OpenStudy (anonymous):

completing the square for the expression 3x^2-6x-5 yeilds...? im really confused

OpenStudy (saifoo.khan):

use another method. that method sucks.

OpenStudy (anonymous):

ik but it says i have to use it

OpenStudy (mimi_x3):

Yeah true the method sucks, its time consuming xD

OpenStudy (anonymous):

3(x^2-2x)-5=0 3(x^2-2x+1)-3 -5=0 3(x-1)^2 -8 = 0 3(x-1)^2 = 8 (x-1)^2 = 8/3 x-1 = +- sqrt(8/3) x=1 +- sqrt(8/3)

OpenStudy (anonymous):

thnx got the answer =]

OpenStudy (anonymous):

it was 3(x-3)^2-14

OpenStudy (mimi_x3):

Which ones right ?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3x^2-6x-5

OpenStudy (anonymous):

why mimi ? I didnt factor the 5. also , wolfram gets the same answer I got.

OpenStudy (mimi_x3):

nvm, i made a stupid mistake xD